Atsie (atsie):

I need some step by step help on this math problem. I know I am to get ride of the fractions somehow, but I really don't know how to start or finish. (pic below)

Atsie (atsie):



OpenStudy (aravindg):

Okay. For a the basic idea is to multiply both sides by the common multiple of denomiNators on LHS and RHS.

OpenStudy (aravindg):

ie we have 1/2 on left, 1/3 on right. Can you think of a multiple of both 3 and 2?

Atsie (atsie):

Wait, what is LHS and RHS??

OpenStudy (texaschic101):

left hand side and right hand side

Atsie (atsie):

ooh, duh! I'm a dunce. So, as for a multiple of both 3 and 2, would it be 4? I had also heard that I would need to subtract these fractions by something, and my mom wrote this on a peice of paper in regards to this problem : (x-4) = 1/3x + 1 1/2 I don't understand what that did though.

OpenStudy (aravindg):

Not 4.

OpenStudy (texaschic101):

common denominator of 3 and 2 is what

OpenStudy (texaschic101):

oops...of 1/2 and 1/3

Atsie (atsie):

Would it be 1? Because 1 is the lowest and is repeated twice?

OpenStudy (texaschic101):

the number that 3 and 2 will both go into evenly is what ?

OpenStudy (aravindg):

Let's try some examples. We cant try to find common multiples for 5 and 4. This means 5* x(some number)=4*y(another number) Clearly 20 is one of the values. 5*4=4*5=20(which is a multiple of both 5 and 4 Similarly for 2 and 5 its 10. 2*5 =5*2=10(which is a multiple of 5 and 2) Note that I am only taking least common multiple here. Now tell me the case for 2 and 3.

OpenStudy (texaschic101):

2,4,6,8,10 3,6,9,12,15 do u see a number that both 3 and 2 go into ?

Atsie (atsie):

Six!!

Atsie (atsie):

Right?

OpenStudy (aravindg):

Finally!

OpenStudy (aravindg):

Once you find that it's just a wave of your magic wand. Multiply both sides by 6 and see what happens.

OpenStudy (texaschic101):

lol..bye bye fractions

Atsie (atsie):

6 x 2 = 12 and 6 x 4 = 24 Both sides by 6 ^^

OpenStudy (texaschic101):

u multiply 1/2 and 1/3 by 6

Atsie (atsie):

oh... 1/2 x 6 is 3 I believe, and 1/3 by 6 is 1.9 Right?

OpenStudy (texaschic101):

both sides by 6

OpenStudy (aravindg):

6*(1/2)*(x-4)=6*(x/3+2)

Atsie (atsie):

What is x though? I don't understand what it resembles.

OpenStudy (aravindg):

x is the variable in your question I just wrote (1/3)x as x/3

Atsie (atsie):

So, 6 * 1/2 is 3 which I then subtract by 4? How on earth do I incorporate x which has no resemblance to any number but just stands there as a letter?

OpenStudy (texaschic101):

6 * 1/2 * (x- 4) = 6 * 1/3x + 6 * 2

Atsie (atsie):

I know, you just repeated what Arav posted up above, but I'm halting the equation when I'm trying to do that. 6 * 1/2 * ??? It say's x, but how do you multiply by a variable that doesn't resemble anything?

OpenStudy (texaschic101):

1/3x is the same as x/3..... 6 * 1/2 is 3...so the lhs will be 3(x - 4).....can u understand that ?

OpenStudy (texaschic101):

rhs.... 6 * (1/3) x + 6 * 2 2x + 12....understand that ?

OpenStudy (texaschic101):

so now u have... 3(x- 4) = 2x + 12 u understand ?

Atsie (atsie):

Okay, so 3 - 4 will be -1 and 2 + 12 will be 14.

OpenStudy (texaschic101):

no... 3(x - 4)....distribute....take the 3 and multiply it by everything in the parenthesis... 3 * x + 3 * - 4 = 3x + (- 12) = 3x - 12....understand so far ?

OpenStudy (texaschic101):

now we have... 3x - 12 = 2x + 12

OpenStudy (texaschic101):

3x - 12 = 2x + 12 3x - 2x = 12 + 12 x = 24

Atsie (atsie):

3 times x You cannot multiply 3 times a letter. Thats what I don't get.

OpenStudy (aravindg):

Its called a variable nt letter in mathematics. You can multiply a variable to a number because variable gets values which ate numbers.

Atsie (atsie):

But how does 3 times x times - 4 = 3x ? What is that x DOING to make the whole equation equal 3x ?

OpenStudy (texaschic101):

you can multiply 3 times x....it is 3x. the x just happens to be a number u dont know yet....so u have to find what that number is....u have to find x

OpenStudy (texaschic101):

3(x - 4)...to get rid of parenthesis, u multiply 3 by everything in the parenthesis... 3 * x + 3 * - 4 = 3x + (- 12) = 3x - 12

Atsie (atsie):

Alright, I think I have an idea of this now. So for part b of the section, how would I write a word problem out?

Atsie (atsie):

@radar Do you think you could maybe look this over and help with part b of my question? I'd appreciate it immensely sir!! ^.^